OpenStudy (umarsback):

Write an equation of a line in slope intercept form that is perpendicular to y = -4x -2 and passes through the point (-16, -11).

OpenStudy (anonymous):

oh boy.

OpenStudy (umarsback):

help please :/

OpenStudy (anonymous):

The first thing that we need to do is to solve 4x+y+2=0 for y. In this example we have: y=−4x−2 We can see that the slope of given line is m=−4. The slope of the perpendicular line is the opposite reciprocal.This means that the perpendicular slope ( m1) can be found by using the formula: m1=−1m. In this case we have: m1=−1m=−1−4=14 We can conclude that the slope of the line we need to find is m=14. Also, we know that the unknown line is passing through point A = ( −16 , −11 ). Now we can use the point-slope form to find the line. The point-slope form is: y−y0=m⋅(x−x0), where m is the slope and (x0,y0) is the given point In our case: m=14, x0=−16, y0=−11 After substitution we have: y−y0y−(−11)yy=m⋅(x−x0)=14⋅(x−(−16))=14⋅x+4−11=14x−7

OpenStudy (anonymous):

ah vvhat it didn't post right lame

OpenStudy (anonymous):

y = 1/4x - 7

OpenStudy (anonymous):

give me one second ill fix up the explanation

OpenStudy (umarsback):

imma read all that xD

OpenStudy (umarsback):

okay

OpenStudy (anonymous):

http://cl.ly/image/1c202Y1S101m

OpenStudy (umarsback):

so y = 1/4x - 7 is the answer?

OpenStudy (anonymous):

yes sir.

OpenStudy (umarsback):

what did you use to solve them .-.

OpenStudy (anonymous):

its in the picture i sent o.o

OpenStudy (umarsback):

i mean what program haha

OpenStudy (umarsback):

did you type all that

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

i used a calculator and vvord

OpenStudy (anonymous):

http://cl.ly/image/0p2S2r1x2B1q

OpenStudy (umarsback):

thanks bro